OpenStudy (anonymous):

You have 332 feet of fencing to enclose a rectangular region. What is the maximum area? A. 6889 square feet B. 6885 square feet C. 110,224 square feet D. 27,556 square feet

OpenStudy (anonymous):

i calculated it and got A as an answer

OpenStudy (jack1):

yep, A

OpenStudy (jack1):

2L + 2W = 332 and L x W = y so from eqn 1, L = 332/2 + W so (332/2 + W) x W = y 166W + W^2 = y when derivative of y = 0 = maximum area y' = 2W + 166 0 = 2W +166 -166 = 2W W = 83 therefore as 2L +2W =332 and w = 83, L must =83 also so L x W =A 83 x 83 = 6889 A

OpenStudy (anonymous):

http://prntscr.com/15hjiz Is its answer B

OpenStudy (anonymous):

since a fence can be a cicle 332 is circumference C=piD=2pir=332 r=332/2pi now Area=pir^2 we have r find the AREA

OpenStudy (anonymous):

http://prntscr.com/15hjiz Is its answer B

OpenStudy (jack1):

@pasta yeah but the q says "rectangular area"